Joe has $6.75 in dimes and quarters. He has 8 more quarters. How many quarters does joe have?
Answer by stanbon(75887)

Joe has $6.75 in dimes and quarters. He has 8 more quarters than dimes. How many quarters does joe have?
-----
Quantity:: q - d = 8
Value:: 25q + 10d = 675 cents
--------------------------------------
Modify for elimination:
5q - 5d = 40
5q + 2d = 135
---------------------
Subtract and solve for "d":
7d = 95
d = 95/7 which is not a whole number.
---------------------
Please check your post.
Cheers,
Stan H.
--------------